Shift Summary for Eve, Feb. 1, 2000


Subject: Shift Summary for Eve, Feb. 1, 2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:49:24 CST


              Shift Summary for eve shift, 01/Feb/00

On Shift: Tschirhart, Bellavance + Wah & Kessler

Hours of beam delivered at all intensities: 6.5/8.5
Hours of KAMI data: 6.3
Accounting of lost time: TeV problems, run change
========================================================================
Runs ended this shift:
  Run 15735 - KAMI_MINBIAS
  Run 15736 - KAMI_KE3 (messed up - not usable)
  Run 15737 - KAMI_KE3 waiting to start

Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: none
========================================================================
Overview of shift:
  Continuation of KAMI data collection.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
              15:30 Take over from day shift.
15735 159 16:14 Beam disappears. Ch.13:"Problem with TeV
                        extraction. Estimate time to beam: 25 minutes."
              16:40 Ch.13:"Problem with TeV extraction. Est. 10
                        minutes."
              16:44 Rick Ford calls to remind us that beam will
                        disappear at 6am tomorrow morning for targeting
                        adjustments.
       183 16:47 Beam returns.
       187 16:53 Beam gone again. Ch.13:"Problem with TeV
                        extraction persist. No estimate."
       191 16:59 Beam returns.
              19:15 Dinner run by Angela.
              21:00 Wilson hall run by Bob.
       406 21:47 Kessler is bored. Wants to do something
                        different. Stop the run. Switch from 20" Be to
                        0" Be.
15736 1 22:01 Start a KAMI_KE3 run.
       5 Remove prescale of 10 from the trigger.
       12 Turned off something that shouldn't have been on
                        anyway (ET boards I think). See log book for
                        this and other changes made en-run.
       32 22:44 Beam gone. Ch.13:"Booster RF problems. Estimate in
                        15 minutes." Stop run (not useful).
              23:26 Ch.13:"Booster RF station 18 is holding off
                        beam. Experts enroute. Estimate 1 hour to beam."
              23:50 Gas checklist done.
              24:00 Several KAMI people are having an impromptu
                        meeting in the control room to decide what to do
                        next. We'll let them tell the next shift what to
                        do.

========================================================================

_______
Angela Bellavance work: (630) 840-2801
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
                                - Arthur C. Clarke



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:49:26 CST